# Protecting Deep Learning Model Copyrights with Adversarial Example-Free Reuse Detection

### v1

@@ VII-A DNN Watermarking and Fingerprinting
Watermarking and fingerprinting are two main approaches for passive DNN intellectual property protection [44]. Watermarking methods embed owner-specific information into DNN for IP violation detection. White-box watermarking methods [13, 14, 15] exploit the over-parameterization property of neural networks to plant secret signatures into the model weights through regularization during training. The ownership can be checked by extracting the signatures from the model weights. Black-box watermarking methods [16, 17] use a backdoor attack to implant a watermark in the model during training. The watermark can be extracted by querying the model with triggering samples. Either way, watermarking methods need to be applied in the training process, which will affect the model performance, and there are also methods [18, 19] proposed for removing the watermarks.
Fingerprinting and testing-based similarity comparison methods are post hoc detection approaches that do not tamper with the training process. They extract characterizing features (sometimes called fingerprints) from the models using carefully generated input samples close to the decision boundary [20] or conferrable adversarial examples [21]. A recent study [22] utilizes universal adversarial perturbations [45] to achieve similar effects. ModelDiff [25] and DeepJudge [24] are the two state-of-the-art testing-based methods. ModelDiff can handle the transfer learning reuse case, and DeepJudge has a white-box detection procedure and achieves a higher detection accuracy on homogeneous reuse cases.
@@ VII-B Active DNN Copyright Protection
Different from passive protection methods, active protection aims to prevent copyright infringement from occurring. For distributed model copies, DNN authentication techniques [46, 47, 48] manage the authorized usage of the source model by encoding the model with an authentication key, so that only authorized users with valid keys can access the model normally, while adversaries with invalid keys can only use the model with greatly reduced performance, thus preventing unauthorized model usage. As for MLaaS mode, model extraction is the major threat, and inference perturbation techniques [49, 50, 51] are proposed to defend the source model by modifying the inference results before returning them to the users.
@@ VIII Conclusion
In this paper, we propose NFARD to protect deep neural network copyright based on measuring the neuron functionality similarity of two models. We devise a linear transformation method which enables NFARD to detect heterogeneous reuse cases that existing approaches could not handle well. To better evaluate copyright protection methods, we construct a reuse detection benchmark covering various commonly used reuse methods and popular model architectures. Empirical evaluation results show the effectiveness and efficiency of our approach in both black-box and white-box settings, outperforming existing methods. In the future, we will explore the application of our approach to other learning tasks and model structures, e.g., language models for text classification.

### v2

@@ VI-A Threats to Validity
VI-B Comparison with Existing Approaches
Here we provide a detailed description of the two state-of-the-art methods, i.e., DeepJudge and ModelDiff, and compare them with NFARD.
DeepJudge first generates a set of adversarial examples for the victim model, then uses these adversarial examples to evaluate the similarity between the victim model and the suspect model through a set of distance metrics, e.g., the difference between the numbers of correctly classified adversarial examples by the two models (referred to as robustness distance). Subsequently, the final decision is determined by comparing the similarity scores against thresholds derived from reference models. The choice of distance metrics for measuring similarity varies between the white-box and black-box settings, as certain metrics rely on the hidden layer outputs of the victim model. Nevertheless, the rationale behind DeepJudge is to exploit the transferability of adversarial examples to detect reuse relationships, which is fundamentally different from our approach. Therefore, DeepJudge requires gradients of the victim model to efficiently generate adversarial examples. It also needs ground truth labels to compute the distance metrics. In addition, DeepJudge cannot handle surrogate models obtained by transfer learning under the black-box setting due to the change of classification task.
Modeldiff follows a similar pipeline to DeepJudge, but differs in the adversarial example generation process and the similarity metrics. ModelDiff generates adversarial examples by maximizing a customized score function, which is designed to ensure a high divergence between the original outputs and the adversarial outputs, and at the same time to increase their diversity. The similarity metric in ModelDiff is the cosine similarity between the two models’ decision distance vectors. A decision distance vector (DDV) of a model $f$ is a vector composed of cosine distances between $f$’s outputs on normal samples and adversarial samples. Therefore, ModelDiff also requires gradients to generate adversarial examples. However, it does not need ground truth labels since its similarity metric is task-agnostic, and it does not differentiate between the white-box and black-box settings. ModelDiff can handle heterogeneous suspect models, but it struggles to effectively detect knowledge distillation and model extraction cases.
In contrast, NFARD is principled on the neural functionality differences among independently trained models, thus it does not require adversarial examples. This feature makes it more efficient and applicable to fully black-box settings where the victim model’s gradients are not accessible. Moreover, NFARD is designed to handle various heterogeneous reuse types, and the experimental results show that it performs better at handling heterogeneous reuse cases than existing methods. The no need for ground truth labels is another minor advantage of NFARD, which makes it more general and applicable to various tasks.
The proposed method still has several limitations. Firstly, although it is applicable to both black-box and white-box settings, the detection accuracy in the black-box setting is lower than that in the white-box setting. This is because lacking the victim model’s internal information makes it more challenging to detect reuse relationships. Secondly, the selection of reference models might pose bias to the detection results. To mitigate this, one could use more reference models, but this would result in a higher computational cost, especially when the victim model is large. Thirdly, our method requires a representative dataset for model characterization, which may not be available, or the number of samples may be limited.
VI-C Generalization to Other Scenarios
The black-box and white-box settings are two common scenarios in DNN reuse detection, in which cases our proposed method has shown good performance. Its applicability to other scenarios, such as distributed learning, remains to be explored. Nevertheless, NFARD is the first method (to the best of our knowledge) that can operate without access to the internals of the victim model, which is a significant step towards practical reuse detection.
We considered seven commonly used reuse techniques in this work, but there are other model modification techniques that could be used to derive surrogate models, such as sparse coding. While these techniques are not covered in this work, the proposed method should remain effective in detecting reuse relationships in these cases, as long as the modified model is intended to retain the victim model’s functionality. In fact, the better the surrogate model preserves the victim model’s behavior, the easier it is for NFARD to detect the reuse relationship.
@@ VII-A DNN Watermarking and Fingerprinting
DNN intellectual property protection can be divided into two categories: active protection and passive protection. Active protection prevents IP infringement by managing the authorized usage or perturbing the inference results, while passive protection aims to enable IP violation detection. Here we focus on passive protection, which includes DNN watermarking and DNN fingerprinting.
Watermarking methods embed owner-specific information into DNN for IP violation detection [44]. White-box watermarking methods [13, 14, 15] exploit the over-parameterization property of neural networks to plant secret signatures into the model weights through regularization during training. The ownership can be checked by extracting the signatures from the model weights. Black-box watermarking methods [16, 17] use a backdoor attack to implant a watermark in the model during training. The watermark can be extracted by querying the model with triggering samples. Either way, watermarking methods need to be applied in the training process, which will affect the model performance, and there are also methods [18, 19] proposed for removing the watermarks.
Fingerprinting methods are post hoc detection techniques that do not tamper with the training process. They extract characterizing features (called fingerprints) from the models using carefully crafted input samples close to the decision boundary [20, 21]. A recent study [22] utilizes universal adversarial perturbations [45] to achieve similar effects. ModelDiff [24] and DeepJudge [23] are the two state-of-the-art fingerprinting methods. ModelDiff can handle the transfer learning reuse case, and DeepJudge has a white-box detection procedure and achieves a higher detection accuracy on homogeneous reuse cases. Besides, [46] proposes a fingerprinting method for protecting copyrights of federated learning models.
@@ VII-B Active DNN Copyright Protection
Different from passive protection methods, active protection aims to prevent copyright infringement from occurring. For distributed model copies, DNN authentication techniques [47, 48, 49] manage the authorized usage of the source model by encoding the model with an authentication key, so that only authorized users with valid keys can access the model normally, while adversaries with invalid keys can only use the model with greatly degraded performance, thus preventing unauthorized model usage. As for MLaaS mode, model extraction is the major threat, and inference perturbation techniques [50, 51, 52] protect the source model by modifying the inference results before returning them to the user.
@@ VIII Conclusion
In this paper, we presented NFARD, a novel deep neural network copyright protection method centered on neuron functionality analysis. The core innovation lies in circumventing the need for adversarial examples, enabling robust model reuse detection using only normal test samples. This makes NFARD particularly well-suited for practical black-box deployments. Additionally, our proposed linear transformation technique effectively tackles heterogeneous reuse scenarios across different architectures and taskss, significantly extending its application scope. Evaluation results demonstrate NFARD’s effectiveness and efficiency, offering a versatile tool for detecting model reuse and safeguarding intellectual property rights against unauthorized replication.
Acknowledgment
This research was sponsored by National Natural Science Foundation of China under Grant No. 62172019 and National Key R&D Program of China under Grant 2022YFB2702200.